An ultrasonic transducer used for medical diagnosis oscillates at a frequency of 6.7MHz=6.7×10⁶Hz. How much time does each oscillation take and what is the angular frequency of the oscillation
1
Expert's answer
2021-08-13T10:20:32-0400

The period of oscillation

"T=\frac{1}{f}=\frac{1}{6.7*10^{8}\:\rm Hz}=1.5*10^{-9}\:\rm s"

The angular frequency of the oscillation

"\omega=2\pi f=2\pi*6.7*10^8=4.2*10^9\:\rm rad/s"